Caspian Sea Drinks' is financed with 68.00% equity and the remainder in debt. They have 10.00-year, semi-annual pay, 5.04% coupon bonds which sell for 97.81% of par. Their stock currently has a market value of $24.49 and Mr. Bensen believes the market estimates that dividends will grow at 3.78% forever. Next year’s dividend is projected to be $2.67. Assuming a marginal tax rate of 23.00%, what is their WACC (weighted average cost of capital)?
Answer #1
Weight of equity = 1-D/A
Weight of equity = 1-0.32
W(E)=0.68
Weight of debt = D/A
Weight of debt = 0.32
W(D)=0.32
Cost of equity
As per DDM
Price= Dividend in 1 year/(cost of equity - growth rate)
24.49 = 2.67/ (Cost of equity - 0.0378)
Cost of equity% = 14.68
Cost of debt
                  K = Nx2
Bond Price =∑ [(Semi Annual Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                  K =10x2
978.1 =∑ [(5.04*1000/200)/(1 + YTM/200)^k]     +   1000/(1 + YTM/200)^10x2
                   k=1
YTM = 5.3252992784
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 5.3252992784*(1-0.23)
= 4.100480444368
WACC=after tax cost of debt*W(D)+cost of equity*W(E)
WACC=4.1*0.32+14.68*0.68
WACC =11.29%